OpenStudy (anonymous):

What is 5-6 divided by 8+7?

Mehek (mehek14):

\(\Large\dfrac{5-6}{8+7}\) like this?

OpenStudy (anonymous):

is it written like 5 - (6 / 8) + 7 ?

OpenStudy (anonymous):

Yes

Mehek (mehek14):

which one?

OpenStudy (anonymous):

And then simplified

OpenStudy (anonymous):

mine or mehek?

OpenStudy (anonymous):

Yours Mehkek14

OpenStudy (worlddomination):

yep

Mehek (mehek14):

\(\Large\dfrac{5-6}{8+7}?\)

Mehek (mehek14):

you have to first subtract 5-6 and add 8+7

OpenStudy (worlddomination):

it would be 15 divided by 1

OpenStudy (i_love_my_nieces):

11.25

OpenStudy (i_love_my_nieces):

but idk if it right

OpenStudy (worlddomination):

no it is not it is 15 because 8+7 = 15 and 6-5 = 1 then you divide them and get 15

Mehek (mehek14):

\(\large\dfrac{5-6}{8+7}=\dfrac{-1}{15}\)

OpenStudy (worlddomination):

so i was partially right

OpenStudy (worlddomination):

forgot abt the negative